Foreverangel:

f g = <27,42> and h = <-12,12> find ⅓ g. <-4, 4> <81, 126> <9 , 14> None of the above

JuniorRozier:

1/3 g = <9,14> bc f(g) = <27,42> divide the variables of g by 3 and bam there you go i believe that math is correct i did my best

Foreverangel:

So is it that you just multiply <27,42> by 1/3?

JuniorRozier:

you divide <27,42> by 3 27/3 = 9 42/3 = 14

JuniorRozier:

or yeahhhhh multiply by 1/3 that works too:)

Foreverangel:

Thank you!

JuniorRozier:

anytime